you are working at a company that manufactures electrical wire. Gold is the most ductile of all metals: it can be streched into incredibly long, thin wires. the company has developed a new technique that will stetch 1.00 g of gold into a wire of length L=2.40 km amd uniform diameter.your supervisor gives you the task of determining the resistance of such a wire at 20.0 degrees celcius.
1
Expert's answer
2020-04-06T09:28:14-0400
"R=\rho^{res}\frac{l}{A}=\rho^{res}\frac{l^2}{V}=\rho^{res}\rho^{den}\frac{l^2}{m}"

"R=(2.44\cdot10^{-8})(19320)\frac{2400^2}{0.001}=2.72\cdot10^{6}\Omega"